Consider a market for avocados with the following demand and supply functions: Pa 150 30a market. (2 points) Suppose that prices are at S40. What is happening in the market? If there is a shortage or excess supply, calculate how much. Display this on your graph created in part 1. (2 points) 6. points) rchased new hydraulic machines to harvest from trees. Assuming that the new supply is given by B = 2Q, + 5, what is the new equilibrium price and quantity? (2 points) Consider two changes in the market related to avocados. First, consumers view avocados and tortilla chips as complements and the price of tortilla chips increases. Second, a new avocado variety is invented that increases the yield per tree. What happens to demand, supply, equilibrium price and equilibrium quantity? Describe each effect qualitatively. Use a graph to help explain your reasoning. (4 points) 9.
Answer #1

Ans 5 Graph of Supply and Demand

when Pd = 150 - 3Qd Equation (1)

Ps = 25 + 2Qs Equation (2)

Equation (1)

Pd 30 60
Qd 40 30

at Pd =30

30= 150-3Qd

-120=-3Qd

40= Qd

at Pd =60

60= 150-3Qd

-90=-3Qd

30= Qd

Equation (2)

Ps 35 45
Qs 5 10

When Ps = 35

35= 25 + 2Qs

10= 2Qs

5= Qs

When Ps = 45

45= 25 + 2Qs

20= 2Qs

10= Qs

3 ㄟLine AB shows Supply and Line CD shows Demand and point E shows the Equilibrium Price where demand is equal to supply.

6) If Price = $ 40

As the price is lower than the Equilibrium price therefore Demand is more than price

7) Excess supply and excess demand will not exist in long run because in case of excess supply then demand then the price of the goods started to decrease which leads to decrease in supply until the supply reaches to demand which is at equilibrium price and in case of Shortage of Supply the demand is high therefore price started to increase until the supply reaches to demand at the equilibrium price.

8)New Ps = 2Qs + 5 Equation 3

compare equation 1 and 3

At equilibrium price the Qs = Qd and Ps = Pd

2Q + 5 =   150 - 3Q

2Q + 3Q = 150 - 5

Q = 145/5

Q = 29

P = (2*29) + 5

P = 58 + 5

= $ 63

So Equilibrium price and quantity is $ 63 and 29 units

9) a) If Avocados and tortilla chips are complementary to each other and there is increase in the price of the tortilla chips then there is decrease in demand of avocados, but there is no such change in the supply pattern , As a result the Equilibrium price moves downwards, as shown in the diagram by Line FG and E2 where price is P1 and Quantity is Q1.

b) If there is new avocado variety is invented that leads to increase in yield per tree then in that case the the supply of the product is increases but there is no such change in the demand pattern,  therefore the Equilibrium price moves downwards as more supply is made at lower price as shown in the diagram by line HI and E3 where price is P2 and Quantity is Q2.

ㄟ Quownh